### v1

# Curriculum Reinforcement Learning from Easy to Hard Tasks Improves LLM Reasoning

Abstract
We aim to improve the reasoning capabilities of language models via reinforcement learning (RL). Recent RL post-trained models like DeepSeek-R1 have demonstrated reasoning abilities on mathematical and coding tasks. However, prior studies suggest that using RL alone to improve reasoning on inherently difficult tasks is less effective. Here, we draw inspiration from curriculum learning and propose to schedule tasks from easy to hard (E2H), allowing LLMs to build reasoning skills gradually. Our method is termed E2H Reasoner. Empirically, we observe that, although easy tasks are important initially, fading them out through appropriate scheduling is essential in preventing overfitting. Theoretically, we establish convergence guarantees for E2H Reasoner within an approximate policy iteration framework. We derive finite-sample complexity bounds and show that, when tasks are appropriately decomposed and conditioned, learning through curriculum stages requires fewer total samples than direct learning. Experiments across multiple domains show that E2H Reasoner significantly improves the reasoning ability of small LLMs (1.5B to 3B), which otherwise struggle when trained with vanilla RL alone, highlighting the effectiveness of our method.
1 Introduction
Large Language Models (LLMs) have demonstrated reasoning capabilities in tasks such as multi-step arithmetic, symbolic planning and code generation. However, the notion of reasoning in LLMs remains vague, with some works equating it to generating intermediate steps during problem solving (wei2022chain, ; cobbe2021gsm8k, ). This view, although intuitive, blurs the line between genuine reasoning and surface-level pattern recognition (stechly2024chain, ; valmeekam2023planning, ). Therefore, we adopt a functional view focusing on generalization, defining reasoning as the ability to extract principles from simpler tasks and apply them to harder ones. Supporting this capability requires training methods that go beyond imitation and help models learn underlying problem-solving strategies.
In this direction, the success of DeepSeek R1 (guo2025deepseek, ) and OpenAI o1 (jaech2024openai, ) shows that reinforcement learning (RL) based post-training enhances reasoning. RL uses task-specific rewards based on output correctness, unlike supervised fine-tuning (SFT), which trains models to imitate fixed input-output examples (zelikman2022star, ). However, RL struggles on harder tasks on which pre-trained models have low zero shot performance (deepseek-math, ; simplerlzoo, ) and since rewards are granted only for correct answers, resulting in sparse learning signals.
Motivated by these limitations, we introduce the Easy 2 Hard Reasoner (E2H), a CRL approach for LLM post-training. Using E2H, we show LLMs can learn tasks they initially failed in the zero-shot setting, challenging the belief reinforced by pass@k evaluations that base models outperform RL-trained ones at a higher $k$ (deepseek-math, ; yue2025does, ) (Fig. 1). We further show that LLaMA models and small models like Qwen 1.5B, often considered incapable of reasoning due to their pre-training data (gandhi2025cognitive, ) or insufficient scale (guo2025deepseek, ), can reason effectively when post-trained with E2H (Fig. 1). This capability comes from using a probabilistic scheduler in E2H that gradually shifts focus from easy to hard tasks, allowing LLMs to build foundational skills on simpler tasks and progressively solve harder ones, leading to improved generalization. On the theoretical front, we provide a comprehensive analysis of curriculum reinforcement learning through the lens of Approximate Policy Iteration. We establish the convergence guarantee for final performance gap and finite-sample analysis for sample complexity bound. Importantly, we prove a sample efficiency comparison that CRL requires fewer total samples than direct learning under appropriate design and learning of curriculum stages, which aligns with our empirical observations. Empirically, we demonstrate that E2H achieves state-of-the-art performance across five reasoning tasks, including planning tasks such as Blocksworld valmeekam2024llms and Countdown (gandhi2024stream, ), and three arithmetic reasoning tasks ((hendrycksmath2021, )(agarwal2021theory, ), and (cobbe2021gsm8k, )), outperforming RL and curriculum baselines.
(a) Pass@$k$ on Countdown
(b) Pass@$k$ on Blocksworld
(c) LLaMA-3.2 reasoning
Figure 1: (a, b) Reinforcement learning (RL) based post-training is commonly believed to primarily improve accuracy at low $k$ values in pass@$k$ evaluation (guo2025deepseek, ; yue2025does, ), we show that our method, E2H Reasoner, a curriculum-based RL (CRL) approach, enables LLMs to solve tasks they previously could not, outperforming base models even at higher $k$. (c) Prior work (gandhi2025cognitive, ) suggests otherwise, we show that LLaMA 3.2 3B can learn to reason effectively with E2H Reasoner.
2 Background and Related Work
Reasoning in Large Language Models remains loosely defined, with interpretations varying by tasks and contexts reasoningsuvery2023 . Prior work describes it as generating logical chains of thought wei2022chain ; wang2022self , performing multi-step deductions (saparov2023language, ; yao2023tree, ; ling2025complex, ), or simulating human-like problem solving (parashar2025inference, ). These views, lack clear boundaries between reasoning and pattern recognition. To address this gap, we are inspired by prior work that frames reasoning as generalization or abstraction (valmeekam2024llms, ; valmeekam2023planning, ), building on the idea that reasoning involves learning core principles and applying them broadly (stechly2024chain, ; huang2024large, ).
Post-training of Large Language Models has emerged as a popular approach to improve reasoning (snell2025scaling, ). These methods are grouped into Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) based techniques. In SFT, the model is trained to imitate outputs from carefully curated human-like reasoning examples (zelikman2022star, ; muennighoff2025s1, ). However, studies have shown that SFT can lead models to overfit to surface-level patterns (chu2025sft, ), limiting generalization. In contrast, RL-based post-training uses task-specific rewards and updates the model through policy optimization algorithms (ppo, ; dpo, ; deepseek-math, ), instead of imitation. This approach has shown greater potential in improving reasoning performance, as demonstrated by the success of models fine-tuned with RL (guo2025deepseek, ; jaech2024openai, ). Still, for inherently difficult tasks that LLMs struggle to solve in zero-shot settings, post-training with RL alone has been insufficient deepseek-math ; simplerlzoo .
Curriculum Learning is a training paradigm that organizes tasks in increasing order of difficulty to facilitate smoother and more effective learning (bengio2009curriculum, ; graves2017automated, ). In the context of RL, it has been applied to help agents acquire complex behaviors by first mastering simpler tasks (narvekar2020curriculum, ). Despite its potential, the application of curriculum-based RL to LLM post-training remains underexplored. Recently efforts have been made to investigate how curriculum-based RL can enhance reasoning and generalization in LLMs (qiu2025wisdom, ; bae2025online, ; simplerlzoo, ). To support gradual learning, some works generate Chain-of-Thought (CoT) data spanning a range of difficulty levels qiu2025wisdom . Others improve learning by removing examples that are too easy or too hard bae2025online , or by maintaining a balanced mix of task difficulties simplerlzoo . Other recent efforts have implemented manual curricula that switch from easy to hard tasks after a fixed number of training iterations (xie2025logic, ; team2025kimi, ). In contrast, our work schedules tasks probabilistically from easy to hard, improving reasoning and generalization to out-of-distribution tasks.
Figure 2: Task Decomposition of Easy 2 Hard Reasoner (E2H). E2H first decomposes the overall task into levels of increasing difficulty, namely trivial, easy, and medium, to help the LLM acquire core skills. As training progresses, E2H schedules harder tasks accordingly. See Section 3.2 for scheduling details.
3 Method
[…]
3.1 Task decomposition for RL post-training
While SFT provides strong supervision signals, chu2025sft suggest that reasoning and generalization ability are more effectively enhanced through RL-based post-training. However, applying RL techniques similar to DeepSeek-R1-Zero (guo2025deepseek, ) to learn complicated reasoning tasks remains challenging. In this work, we analyze these challenges in two key aspects, the distribution gap and reward design.
Challenge 1: Distribution gap. Learning tasks that exceed the base LLM’s reasoning capabilities introduce significant learning challenges. These challenges are often caused by a non-trivial distribution gap between the model pre-training source data distribution $d_{0}$ and the target data distribution $d_{K}$. As shown in Fig. 2, because rewards are only given for correct outputs, large distribution shifts can lead to low accuracy and sparse reward signals (deepseek-math, ; simplerlzoo, ). Moreover, fitting the model to a single target distribution can lead to overfitting and memorization, undermining the model’s generalization and reasoning ability.
Challenge 2: Reward design. Challenging reasoning tasks often require LLMs to combine multiple skills to arrive at a solution. While designing a fine-grained, step-by-step reward function could potentially guide the model effectively, such design is generally task-specific and labor-intensive. For example, a computer science student is supposed to learn basic mathematics and linear algebra before learning machine learning. Similarly, a typical Countdown (gandhi2024stream, ) task involves skills like basic arithmetic, estimating the distance to the goal, and backtracking. While it is possible to include a supervision signal for each intermediate step, doing so is not sustainable and generalizable across diverse tasks.
[…]
In this work, we simplify a hard reasoning task into 3 progressively challenging tasks, namely, trivial, easy, and medium, then we adapt our pre-trained LLM to these tasks in a sequential curriculum. We adopt the reward design from guo2025deepseek , where outputs with the correct format receive partial rewards, and full rewards are given only when both format and answer are correct. In Section 3.3, we provide a theoretical justification that, under fixed sampling and training resources, learning step by step leads to better performance than training directly on the hard task.
Figure 3: Illustration of cosine scheduling.
3.2 Training schedulers for LLM reasoning ability
While task decomposition simplifies RL post-training, traditional curriculum learning poses two main challenges, mainly task forgetting and overfitting, due to the rigid progression through tasks in a fixed task order.
Challenges: Task forgetting and overfitting. The first challenge, task forgetting, refers to the degradation in performance on earlier (easier) tasks as the model adapts to later (harder) tasks. According to the traditional model generalization literature (arjovsky2019invariant, ; gulrajani2020search, ), the task distribution shifts are considered as explicit signals for the model generalization direction; thus, retaining strong performance across all task distributions is essential for generalization. Therefore, the task forgetting will undermine the model’s generalization capability, i.e., the reasoning ability. Task overfitting, the second challenge, arises when the model overfits to trivial tasks and prefers simplistic patterns or short answers that bypass meaningful reasoning. This phenomenon is called reward hacking (laidlaw2025correlated, ), where the model exploits shortcuts on easy tasks and fails to learn harder ones, resulting in poor reasoning performance. To address both challenges, we explore different training sampling techniques, forming four different scheduling strategies as follows.
Traditional scheduling. We first formulate the traditional sequential curriculum learning sampler with $T$ training steps as $\mathbf{S}_{\text{trad}}(t,k)=1$ when $\tau_{k}\leq t\leq\tau_{k+1}$, otherwise, 0, where $t$ denotes the current training step; for $K$ tasks $k=1,\ldots,K$, $\tau_{k}$ denotes the threshold when the curriculum learning proceeds to the $k$-th stage, while $\tau_{1}=0$ and $\tau_{K+1}=T$. The output of the sampler denotes the probability of sampling data from the $k$-th task; therefore, at the $t$ step, the sampling distribution will be [$\mathbf{S}_{\text{trad}}(t,1)$, $\mathbf{S}_{\text{trad}}(t,2)$, …, $\mathbf{S}_{\text{trad}}(t,K)$].
Balanced scheduling. To avoid forgetting, the simplest way is to mix all data with different difficulties together and sample randomly, which can be considered as a trivial case of curriculum learning. Alternatively, this can be interpreted as the default behavior of any policy optimization algorithm (deepseek-math, ; ppo, ), where training occurs without considering task difficulty. This balanced sampler can be written as: $\mathbf{S}_{\text{balanced}}(t,k)=\frac{1}{K},$ where each task difficulties have the same probability to be selected at each training step. Although this is an efficient way to avoid forgetting, this sampling introduces harder tasks too early, leading to sparser rewards and suboptimal CRL.
Cosine scheduling (E2H-C). To alleviate both the reward sparsity and forgetting issues, we propose a non-parametric scheduling strategy, namely, cosine sampling. This strategy can be written as: $\mathbf{S}_{\text{cosine}}(t,k)=\alpha_{t}\cdot(K-k-1)+(1-\alpha_{t})\cdot k,\,\,\,\text{and}\,\,\,\alpha_{t}=0.5\cdot(1+\operatorname{cos}(\frac{\pi t}{T})),$ where the resulting probabilities need to be normalized before sampling. Intuitively, this cosine sampler sets both the initial and ending sampling probabilities simply according to their ordinal number and interpolates the intermediate probabilities using a cosine function. As shown in Figure 3, the easiest task has the highest probability of being sampled at the beginning, and has the lowest probability of being sampled at the end.
Gaussian scheduling (E2H-G). Although cosine scheduling addresses reward sparsity and forgetting, the parameter-free design limits flexibility in handling issues like trivial task overfitting and fine-grained control over different learning stages. Empirically, while adding the trivial task can boost the model performance, it is also easy for the model to overfit to trivial tasks. To overcome this challenge, we propose a Gaussian scheduling strategy inspired by the Gaussian mixture model (reynolds2009gaussian, ).
Figure 4: Gaussian Sampler. (a) This figure represents the Gaussian sampling process. (bcd) These figures denote the sampling probabilities of different tasks changing along the training steps with different Gaussian sampler hyperparameters.
As shown in Figure 4 (a), in a one-dimensional space, we assume the data distributions of tasks follow Gaussian distributions with the same variance $\sigma$. The means of the adjacent task’s Gaussian distributions are assumed to have the same distance $1$, i.e., $\mu_{k}=k-1$. Then the sampling probability is defined as the likelihood of a given position $x_{t}$ belonging to different task Gaussian distributions, similar to the expectation–maximization algorithm (mclachlan2008algorithm, ). Therefore, the Gaussian scheduling strategy can be expressed as:
$\mathbf{S}_{\text{Gaussian}}(t,k)=\operatorname{exp}\left(-\frac{(x_{t}-\mu_{k})^{2}}{2\sigma^{2}}\right),\ \mbox{and}\ \ x_{t}=\left(\frac{t}{T}\right)^{\beta}\left(K-1\right),$ (1)
where we ignore the normalization term, and the probabilities will be normalized for sampling. In this sampling scheduler, we only have two hyperparameters, i.e., $\sigma$ and $\beta$. While the variance $\sigma$ controls the sampling concentration, $\beta>0$ controls the sampling position $x_{t}$’s moving speed. When $\sigma$ is smaller, the training is more focused, more similar to traditional curriculum learning. When $\beta<1$, the sampling process will assign fewer training steps focusing on easier tasks and train harder tasks longer, avoiding easy task overfitting. As shown in Figure 4 (bcd), we use three typical hyperparameter settings for this Gaussian sampler, i.e., $\beta=0.25\ \sigma=0.75$, $\beta=0.75\ \sigma=0.25$, and $\beta=0.5\ \sigma=0.5$.
3.3 Theoretical Analysis
In this section, we analyze our curriculum reinforcement learning framework for LLMs under the lens of Approximate Policy Iteration (API). We follow the theoretical structure of chen2022approximate ; scherrer2014approximate , adapting it to our curriculum setting. Specifically, we show how sequentially solving interpolated curriculum distributions enables convergence guarantees and improved sample complexity bounds, under function approximation errors. Our analysis considers action-value functions and explicitly tracks the impact of approximation and distribution shift across curriculum stages. All proofs are presented in Appx. A.
3.3.1 Theoretical Setup
Recall our MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$. We denote the policy space as $\Pi$, where each policy $\pi\in\Pi$ is a mapping from states to distributions over actions, $\pi:\mathcal{S}\to\Delta(\mathcal{A})$. In our curriculum setting, we introduce a sequence of MDPs $\{{\mathcal{M}_{k}}\}_{k=1}^{K}$ where each MDP shares the same state and action spaces but may differ in reward functions and/or transition dynamics. Each MDP induces a state visitation distribution $d_{k}$ under the optimal policy for that curriculum, with $\{d_{k}\}_{k=1}^{K}$ interpolating between an easy source distribution $d_{1}$ and the final hard task $d_{K}$. Let $\pi_{k}$ be the learned policy at curriculum step $k$, and let $\pi^{*}_{K}$ be the optimal policy under the final target task. The goal is to learn a sequence of policies $\{\pi_{k}\}_{k=1}^{K}$ such that the final policy $\pi_{K}$ performs well under the target distribution $d_{K}$.
The action-value function of a policy $\pi$ is defined as $Q^{\pi}(s,a):=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\middle|s_{0}=s,a_{0}=a,a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a_{t})\right].$ Let $\mathcal{T}^{\pi}$ be the Bellman operator for policy $\pi$, defined as $\mathcal{T}^{\pi}Q(s,a):=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a),a^{\prime}\sim\pi(\cdot|s^{\prime})}\left[Q(s^{\prime},a^{\prime})\right].$ We define the optimal Bellman operator $\mathcal{T}$, $\mathcal{T}Q(s,a):=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}\left[\max_{a^{\prime}}Q(s^{\prime},a^{\prime})\right].$ The fixed point of the optimal Bellman operator is $\mathcal{T}Q^{*}=Q^{*}$.
The Approximate Policy Iteration (API) framework is a generalization of classical policy iteration that accommodates function approximation and inexact updates. It serves as a foundational tool for analyzing practical reinforcement learning algorithms including actor-critic and deep RL methods. The API algorithm alternates between two stages. At iteration $k$, the algorithm performs two algorithmic steps. $(i)$ Policy Evaluation: Given policy $\pi_{k}$, compute an approximate estimate $\widehat{Q}_{k}$ of its action-value function $Q^{\pi_{k}}$, $(ii)$ Policy Improvement: Update the policy to $\pi_{k+1}$, which is greedy or approximately greedy with respect to $\widehat{Q}_{k}$. In the context of CRL, API provides a natural structure to analyze how sequentially adapted policies over interpolated curriculum distributions evolve toward the optimal policy and influence the final policy $\pi_{K}$.
We adopt the following API assumptions from chen2022approximate ; scherrer2014approximate , adapted to the curriculum setting. Let $Q_{k}:=Q^{\pi_{k}}$ be the action-value function for the policy $\pi_{k}$ at curriculum $k$.
Approximate Policy Evaluation. At each iteration $k$, a function approximator $\widehat{Q}_{k}$ is used to estimate $Q_{k}$. The estimated Q-function $\widehat{Q}^{\pi_{k}}_{k}$ satisfies $\|\widehat{Q}^{\pi_{k}}_{k}-Q^{\pi_{k}}_{k}\|_{\infty}\leq\delta_{k}.$
Approximate Greedy Policy Improvement. Let $\pi_{k+1}$ be an $\epsilon_{k}$-greedy policy with respect to $\widehat{Q}_{k}$: $\mathbb{E}_{s\sim\mu_{k}}\left[Q^{\pi_{k+1}}(s,\pi_{k+1}(s))\right]\geq\mathbb{E}_{s\sim\mu_{k}}\left[\max_{a}\widehat{Q}_{k}(s,a)\right]-\epsilon_{k},$ for a sampling distribution $\mu_{k}$ at step $k$, i.e., $\|\mathcal{T}\widehat{Q}^{\pi_{k}}_{k}-Q^{\pi_{k+1}}_{k}\|_{\infty}\leq\epsilon_{k}.$
Distribution Mismatch (Concentrability). Let $\mu_{k}$ be the sampling distribution for policy improvement at step $k$, and $d_{k}$ the true task distribution. We assume the distribution mismatch between $\mu_{k}$ and $d_{k}$ is bounded by $C_{k}$, i.e., $\sup_{s\in\mathcal{S}}\frac{d_{k}(s)}{\mu_{k}(s)}\leq C_{k}.$
Bounded Curriculum Drift. For the weighted norm, $\|Q\|_{d_{k}}=\sqrt{\mathbb{E}_{s\sim d_{k}}[\max_{a}Q(s,a)^{2}]}$. The deviation between successive optimal Q-functions satisfies $\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}\text{ is bounded for all }k.$
3.3.2 Convergence Guarantee
Let $Q^{*}_{K}$ be the optimal Q-function for the final task under distribution $d_{K}$, and let $Q^{\pi_{K}}$ be the Q-function of the policy learned at the final step. Define the performance loss of the final policy compared to the optimal target policy as: $L_{K}:=\|Q^{*}_{K}-Q^{\pi_{K}}\|_{d_{K}}.$
Theorem 3.1 (CRL Performance Guarantee).
Let $T$ be the number of API policy updates within each task. $\beta>0$ is a tunable parameter for stepsizes specified in chen2022approximate . Under the approximate greedy update and evaluation error assumptions above, the final performance gap $L_{K}$ satisfies:
$\mathcal{E}_{K}\leq\sum_{k=1}^{K}\left(\gamma^{T}\eta_{k}+\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\right)+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}},$
where $\eta_{k}:=\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{\infty}$ is the per-task Bellman error, $\delta_{k}$ is the evaluation error, and $\epsilon_{k}$ is absorbed in $\eta_{k}$.
The first term represents the convergence bias of the actor, and goes to zero geometrically fast as $T$ goes to infinity. The second term captures accumulated evaluation errors, which involves the stochastic error due to sampling and the error due to function approximation The third term captures the error introduced by the policy update rule, which can be made arbitrarily small by using large enough $\beta$. Alternatively, we can use geometrically increasing stepsizes, in which case the third term goes to zero at a geometric rate. The last term captures the deviation in optimal Q-functions across curriculum stages, representing the cumulative gap between intermediate curriculum-optimal values and the final optimal value, which we refer to as the curriculum approximation error. This decomposition highlights the dual effect of CRL in improving sample efficiency (small $\delta_{k},\epsilon_{k}$) and ensuring smooth interpolation (small $\|Q^{*}_{K}-Q^{*}_{k}\|$).
3.3.3 Finite Sample Approximation Error Analyses
In this section, we perform finite-sample analysis. Considering the critic, i.e., how to obtain an estimate of $Q^{\pi}$, we can estimate the Q-function $Q^{\pi}$ of a given target policy $\pi$ using TD-learning. In TD-learning, especially when the state-action space size is large, the use of function approximation is natural. In linear function approximation, a set of basis vectors is selected to approximate the target value function $Q^{\pi}$ using linear combinations of the basis vectors. Let $\Phi\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times d}$ be the matrix of basis vectors $\Phi=[\phi_{1},\cdots,\phi_{d}]$. Then, the goal is to find from the linear subspace $\mathcal{Q}=\{\hat{Q}_{w}=\Phi w\mid w\in\mathbb{R}^{d}\}$ the “best” approximation of the $Q$-function $Q^{\pi}$, where $w\in\mathbb{R}^{d}$ is the weight vector.
Following Section 3 and 4 of chen2022approximate , we derive the finite-sample theorem for CRL. Theorem A.1 and its analyses are provided in Appx. A.1.1. The result of Theorem A.1 suggests that smoother curriculum trajectories and smaller approximation errors under finite samples, which is closely associated with the stepsize and number of updates per iteration, can lead to better final policy performance.
Following chen2022approximate , let $\mathcal{K}_{\mathcal{S}\mathcal{A}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ be a diagonal matrix and let $\mathcal{K}_{\mathcal{S}\mathcal{A},\min}$ be the minimal diagonal entry. $J_{k}$ is the number of critic updates per policy iteration at curriculum $k$. The bootstrapping parameter $n$ is chosen such that $\gamma_{c}:=\gamma^{n}/\sqrt{\mathcal{K}_{\mathcal{S}\mathcal{A},\min}}<1$. $L_{k}$ is the parameter defined by $1+(\gamma\rho_{\text{max},k})^{n}$, $\mathcal{E}_{\text{approx},k}:=\sup_{\pi}\left\|Q^{\pi_{k}}_{c,\rho}-\Phi w^{\pi_{k}}_{c,\rho}\right\|_{\infty}$ is the critic’s function approximation error, $\|Q^{*}_{K}-Q^{*}_{k}\|_{\infty}$ is the gap between the curriculum subtask and the final task, and $\lambda_{\min}$ is the mininum eigenvalue of the positive definite matrix $\Phi_{T}\mathcal{K}_{\mathcal{S}\mathcal{A}}\Phi$. The term $N_{2,1}=\sum_{k=1}^{K}\frac{2\gamma\mathcal{E}_{\text{approx},k}}{(1-\gamma)^{2}}$ in Theorem A.1 is the function approximation error and equals zero when using a complete basis. A term of similar form is present in all existing work studying RL with function approximation (munos2003error, ; agarwal2021theory, ). Based on Theorem A.1, we next derive the sample complexity result.
Theorem 3.2 (Sample Complexity).
For a given accuracy level $\epsilon>0$, to achieve $E[\|Q^{*}_{K}-Q^{\pi_{K}}\|_{\infty}]\leq\epsilon+N_{2,1}+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}},$ the total number of samples (i.e., the integer $T\sum_{k=1}^{K}J_{k}$) required across all $K$ curriculum stages is of the order:
$O\left(\sum_{k=1}^{K}\frac{\log^{3}(1/\epsilon_{k})}{\epsilon_{k}^{2}}\cdot\tilde{O}\left(\frac{L_{k}^{2}n}{(1-\gamma)^{7}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)\right),$
where $\epsilon_{k}$ is the target accuracy for curriculum $k$, and $\sum_{k=1}^{K}\epsilon_{k}\leq\epsilon$.
Let $M_{\text{CRL}}=\sum_{k=1}^{K}M_{k}$ be the total number of samples needed by CRL with $K$ curriculum steps to achieve error $\epsilon$on the final task, and let $M_{\text{Direct}}=m*M_{K}$ be the number of samples needed by direct learning, where $m>1$ represents the relative difficulty factor of direct learning. Under geometric error and $L_{k}$ allocation $\epsilon_{k}=\epsilon_{K}\cdot e^{K-k}$ and $L_{k}=\frac{L_{K}}{l^{K-k}}$ for curriculums, we have:
$M_{\text{CRL}}<M_{\text{Direct}}\iff\frac{(e*l)^{2(1-K)}-1}{1-(e*l)^{2}}<m-1.$ (2)
The geometric error and $L_{k}$ allocation reflect the curriculum gradually increases in difficulty while allowing larger errors in earlier stages. Since the function $f(x)=\frac{x^{1-K}-1}{1-x^{2}}$ monotonically decrease for $x>1$ when $K$ is a integer larger than 1, the final condition can be reasonably satisfied, e.g., with $K=3$, $e*l=1.4$, and $m=1.8$ in practice.
The first half of Theorem 3.2 analysis highlights the dual benefit of curriculum design in CRL. First, by constructing intermediate distributions $d_{1},\dots,d_{K-1}$ close to $d_{K}$, the curriculum error term $\|Q^{*}_{K}-Q^{*}_{k}\|$ can be made small. Second, easier curriculum tasks improve estimation accuracy and yield more stable approximate greedy updates, enhancing sample efficiency and policy improvement. For the second half of Theorem 3.2, the final condition in Eq. 2 is satisfied when curriculums are appropriately learned, the allocation of accuracy targets $\epsilon_{k}$ is gradually optimized across curriculum steps, and the designed curriculums with increasing difficulties effectively bridge the gap between source and target distributions. This mathematical derivation shows CRL can theoretically require fewer total samples than direct learning on the final task, which aligns with experimental observations.
4 Experiments
Table 1: Impact of task decomposition for LLM post-training. Trivial and easy examples help the model learn core principles that enable success on harder tasks.
Blocksworld Countdown MATH
Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD
Hard 0.0 0.0 0.0 0.0 0.0 0.0 43.9 16.4 18.1 6.5 82.3 64.7 53.4 38.2 20.6
Med + Hard 2.0 0.0 0.0 0.0 0.0 12.9 47.8 33.1 19.2 8.8 87.1 68.9 58.1 42.5 21.0
Easy + Med + Hard 0.0 55.5 15.5 0.0 0.0 62.5 79.3 30.1 21.2 9.5 84.8 68.2 56.8 42.5 21.5
Trivial + Easy + Med + Hard 98.0 100 83.3 21.1 2.6 96.1 64.9 28.8 18.1 9.2 87.2 72.0 61.6 46.3 25.7
Table 2: Effect of scheduling strategies in LLM post-training. We compare balanced scheduling, traditional curriculum learning (CL), and our proposed E2H Reasoner variants, namely, E2H-G and E2H-C. CoT is reported as a reference.
Blocksworld Countdown MATH
Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD
CoT 4.0 0.0 0.0 0.0 0.0 16.0 5.6 1.7 0.1 0.1 40.1 27.9 22.7 17.6 8.2
Balanced 98.0 100 84.5 26.3 5.3 96.1 64.9 28.8 18.1 9.2 87.2 72 61.6 46.3 25.7
CL 46.0 100 45.2 5.8 0.7 57.7 85.8 57.2 31.5 12.6 86.2 71.5 62.4 46.7 25.6
E2H-G (0.25, 0.75) 98.0 100.0 95.3 32.9 7.3 98.9 87.3 51.4 18.9 7.3 85.5 72 64.1 47.9 26.5
E2H-G (0.5, 0.5) 100 100 34.5 10.5 0.7 95.2 84.2 48.1 28.1 14.2 85.3 71.7 62.5 48.7 27.6
E2H-G (0.75, 0.25) 98.0 93.3 17.9 2.0 0.0 95.7 56.0 28.8 17.1 10.2 86.0 72.4 62.0 46.7 26.3
E2H-C 100 100 15.5 0.0 0.0 96.7 64.0 25.9 15.8 6.4 84.6 69.6 63.0 47.6 28.6
In practice, we conduct experiments to investigate the following research questions. RQ1: What role does task decomposition play in RL-based post-training? RQ2: How does task scheduling impact the learning process? RQ3: Can small-scale LLMs learn to reason on hard tasks?
4.1 Experimental Setup
We evaluate our method on a diverse set of reasoning and planning tasks, covering both datasets with and without human-annotated difficulty levels. For datasets with difficulty labels, such as Blocksworld (using plan length)(valmeekam2023planning, ), Countdown (using number of operations)(gandhi2025cognitive, ), and MATH (using problem levels) (hendrycksmath2021, ), we use the provided annotations and define an out-of-distribution (OOD) split to assess generalization. For datasets without explicit difficulty, namely GSM8K (cobbe2021gsm8k, ) and AQuA (aqua, ), we estimate difficulty using the zero shot error rate of the base model. Specifically, for each question in the training set, we generate 20 responses and compute the error rate as 1Number of Correct $1-\frac{\text{Number of Correct Responses}}{20}$ and bucket the samples into trivial, easy, medium, and hard, based on quartiles Fig. 4.2. More dataset details are in the Appendix. We conduct experiments using three LLMs spanning 1.5 to 3 billion parameters, namely Qwen 2.5/1.5B Instruct, Qwen 2.5/3B Instruct (yang2024qwen2, ), and LLaMa 3.2 3B Instruct (grattafiori2024llama, ).We use Qwen 1.5B for research questions RQ1 and RQ2.
4.2 Baselines
Firt, we report the Chain of Thought (wei2022chain, ) (CoT) performance for all models. All post-training experiments use GRPO (deepseek-math, ) as the reinforcement learning algorithm (c.f. Appendix for implementation). GRPO, by default, employs balanced scheduling over all tasks, which we use as our baseline and refer to as GRPO in Table 3. We also assess whether models can learn directly from the most challenging examples by training only on the Hard and OOD subsets. In addition, we evaluate traditional curriculum learning (CL), which introduces tasks in a fixed order from easy to hard. Finally, we provide a comparison against SFT in the Appendix.
Table 3: Results of E2H Reasoner across three models on Blocksworld (valmeekam2023planning, ), Countdown (gandhi2024stream, ) and MATH hendrycksmath2021 . Our method consistently improves performance especially on HARD and OOD tasks, demonstrating effective reasoning with small-scale LLMs.
Blocksworld Countdown MATH
Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD
Qwen 1.5B Instruct CoT 4.0 0.0 0.0 0.0 0.0 16.0 5.6 1.7 0.1 0.1 40.1 27.9 22.7 17.6 8.2
GRPO (All) 98.0 100 83.3 21.1 2.6 96.1 64.9 28.8 18.1 9.2 87.2 72.0 61.6 46.3 25.7
GRPO (Hard) 0.0 0.0 0.0 0.0 0.0 0.0 43.9 16.4 18.1 6.5 82.3 64.7 53.4 38.2 20.6
GRPO (OOD) 0.0 0.0 0.0 0.0 0.0 3.1 23.1 18.1 11.3 5.3 37.0 21.1 15.0 8.5 3.7
CL 46.0 100 45.2 5.8 0.7 57.7 85.8 57.2 31.5 12.6 86.2 71.5 62.4 46.7 25.6
E2H-G 98.0 100 95.3 32.9 7.3 95.2 84.1 48.1 28.1 14.2 85.3 71.7 62.5 48.7 27.6
E2H-C 100 100 15.5 0.0 0.0 96.7 64.0 25.9 15.8 6.4 84.6 69.6 63.0 47.6 28.6
Qwen 3B Instruct CoT 1.0 6.7 7.1 1.0 0.0 24.0 15.9 4.4 0.5 0.0 68.0 55.8 48.8 33.4 20.0
GRPO (All) 100 100 88.1 38.3 14.6 99.6 89.3 73.4 41.7 9.4 90.1 78.5 71.9 57.4 36.8
GRPO (Hard) 50.0 42.2 86.9 72.4 48.3 1.6 49.5 38.7 27.7 12.6 90.6 78.4 70.5 58.0 36.0
GRPO (OOD) 0.0 13.3 9.5 0.0 0.0 0.4 25.1 21.2 13.8 8.5 91.3 78.1 70.0 58.7 36.8
CL 100 100 77.3 42.1 20.5 39.0 91.0 72.9 38.7 12.5 90.6 78.9 72.5 57.2 36.6
E2H-G 100 100 96.4 53.3 23.2 98.5 90.8 71.0 43.5 19.4 90.4 81.0 71.5 58.2 37.2
E2H-C 100 100 94.1 52.0 22.5 100 90.4 69.7 35.7 12.6 90.4 80.0 73.4 58.9 38.1
LLaMa 3.2 3B Instruct CoT 24.0 0.0 1.2 1.0 0.0 37.1 4.6 0.3 0.0 0.0 65.9 44.6 35.2 24.1 13.6
GRPO (All) 100 100 94.1 38.9 13.3 99.9 89.5 71.6 47.9 2.7 65.9 47.0 36.0 22.0 10.2
GRPO (Hard) 0.0 0.0 0.0 0.0 0.0 40.5 33.8 3.0 9.7 1.4 22.7 14.4 10.3 7.5 0.3
GRPO (OOD) 0.0 0.0 0.0 0.0 0.0 8.0 0.3 0.0 0.0 0.0 63.6 39.0 31.3 19.0 7.6
CL 100 0.0 0.0 0.0 0.0 17.2 36.0 22.7 11.2 4.1 74.1 54.1 43.9 28.0 12.5
E2H-G 100 100 98.8 44.1 17.6 95.0 89.9 73.3 46.5 24.3 78.7 58.4 46.4 32.3 14.5
E2H-C 100 0.0 0.0 0.0 0.0 100 55.3 0.0 0.0 0.0 74.8 60.6 48.3 34.3 15.8
Table 4: Performance of E2H Reasoner on GSM8K and AQuA, which lack human difficulty labels. We use error rates to create difficulty-based splits of the data, which are then used for scheduling. Fig. 4.2 illustrates these splits.
Qwen 1.5B Instruct
GSM8K AQuA
Trivial Easy Med Hard Avg Trivial Easy Med Hard Avg
CoT 90.2 87.3 76.5 38.1 67.7 70.8 51.3 20.8 2.6 40.9
GRPO 99.0 95.3 84.1 49.9 77.1 95.8 68.0 48.6 21.0 63.3
CL 98.0 97.2 85.8 52.2 78.6 88.8 72.2 36.1 18.4 58.6
E2H-G 97.6 94.7 89.0 51.8 78.7 90.2 81.9 43.0 34.2 66.1
E2H-C 98.0 95.3 83.9 46.6 75.7 86.1 72.2 48.6 26.3 62.5
Figure 5: GSM8K Difficulty distribution based on error rates, grouped into quartiles.
4.3 Experimental Results
We report results addressing the research questions listed above. We examine how task decomposition impacts LLM post-training using Qwen-1.5B-Instruct (RQ1) with a balanced scheduler (Table 1). We find that including trivial and easy examples helps the model build core skills. This enables effective transfer from simpler to harder tasks and better OOD performance, consistent with our view of reasoning as learning core principles and applying them to harder tasks. Practically, easier tasks yield dense rewards, helping the model begin learning effectively and gradually adapt to harder tasks.
Next, we examine how scheduling impacts post-training (RQ2). Balanced scheduling over mixed-difficulty data serves as a strong baseline but lacks structure, leading to suboptimal learning. Traditional curriculum learning (CL) schedules tasks in a fixed order, which can cause forgetting of earlier tasks or overfitting to easier ones. In our experiments, easier tasks help initiate learning due to their dense rewards, but overexposure hinders generalization to harder, sparse-reward tasks. This motivates the need for more effective scheduling and we address these limitations with our proposed E2H-C and E2H-G (see Section 3.2).For tasks like MATH, where models perform reasonably well across all difficulty levels (as seen with CoT in Table 2), the cosine scheduling in E2H-C is effective, beginning with a focus on trivial and easy tasks and gradually shifting toward harder ones. However, on tasks like Blocksworld, where rewards are sparse and harder examples are more challenging, this leads to overfitting and degraded performance. E2H-G addresses this issue by using a Gaussian schedule that quickly decays the sampling probability of trivial and easy tasks. As shown in Fig. 4, it provides enough exposure early on to support initial learning while rapidly shifting focus to harder examples. This prevents overfitting and improves generalization in sparse-reward settings where E2H-C struggles. Further qualitative analyses are provided in the Appendix.
Building on these insights, we investigate whether models can learn to reason directly from difficult examples (RQ3). As shown in Table 3, the answer is largely no. For instance, Qwen-2.5 1.5B when trained on Level 5 MATH directly, underperforms the CoT baseline on the model. Moreover, stronger models like Qwen-2.5 3B can learn Hard Blocksworld tasks directly but fail at the trivial and easy tasks, indicating a lack of basic understanding. These failures highlight the limitations of relying on difficult examples alone. In contrast, E2H Reasoner guides learning by scheduling tasks from easy to hard, improving generalization, as reflected in stronger OOD performance (RQ3). Note that for E2H-G we report the best numbers out of our 3 parameter settings with extensive results in the Appendix. Our results show that small-scale LLMs can learn to reason if taught with task decomposition and appropriate scheduling. Finally, E2H Reasoner remains effective even without human difficulty labels by using error rates as a proxy for difficulty (Table 4).
5 Summary and Limitations
We introduce the E2H Reasoner (E2H), a CRL-based method for LLM post-training. E2H enables models to learn tasks they initially failed at by scheduling tasks from easy to hard. E2H challenges the assumption that small LLMs cannot reason and demonstrates strong empirical performance supported by theoretical analysis, offering convergence guarantees and improved sample efficiency over direct RL. Despite the strengths of our method, it has limitations, we explore only simple, intuitive schedulers, and more adaptive, advantage-based strategies could yield further gains. Additionally, we have not yet scaled post-training to larger models. In summary, E2H provides a scalable, theoretically grounded, and practical method for LLM reasoning.
Acknowledgments
This work was supported in part by National Institutes of Health under grant U01AG070112 and ARPA-H under grant 1AY1AX000053.
Appendix A Proofs
We now present the proofs for the main theoretical results stated above. Our analysis closely follows the derivation style of chen2022approximate ; scherrer2014approximate , extending it to the curriculum setting.
A.1 Proof of Theorem 3.1
Proof.
We aim to derive a tight bound for the final performance error in task $\mathcal{T}_{K}$:
$\mathcal{E}_{K}:=\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}},$
where $Q^{*}_{K}$ is the optimal Q-function for the final task $\mathcal{T}_{K}$, and $Q^{\pi_{K}}_{K}$ is the Q-function under the learned policy $\pi_{K}$.
We begin by analyzing the policy performance error under API. For each task $k$, we define:
$\eta_{k}:=\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{\infty}.$
By norm monotonicity, we have:
$\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}}\leq\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{\infty}.$
We proceed by recalling the result from Theorem 2.1 of chen2022approximate for a fixed task $k$, applied with a step size $\beta$:
$\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{\infty}\leq\gamma^{T}\|Q^{*}_{k}-Q^{\pi_{k}^{0}}_{k}\|_{\infty}+\frac{2\gamma}{1-\gamma}\sum_{t=0}^{T-1}\gamma^{T-1-t}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}},$
where $\delta_{k}$ is the approximate value evaluation error at curriculum step $k$; $\gamma^{T}\|Q^{*}_{k}-Q^{\pi_{k}^{0}}_{k}\|_{\infty}$ reflects initialization at $\pi_{k}^{0}$; and $\beta>0$ is the API soft update step size parameter. The finite geometric series
$\sum_{t=0}^{T-1}\gamma^{T-1-t}$
has closed-form expression
$\sum_{i=0}^{T-1}\gamma^{i}=\frac{1-\gamma^{T}}{1-\gamma}\text{ for }\gamma\neq 1\quad\text{and}\quad\sum_{i=0}^{T-1}\gamma^{i}=1,\text{ for }\gamma=1.$
Thus we simplify with
$\sum_{i=0}^{T-1}\gamma^{i}=\frac{1-\gamma^{T}}{1-\gamma}.$
Since we denote:
$\eta_{k}:=\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{\infty};$
then, the bound becomes:
$\eta_{k}\leq\gamma^{T}\eta_{k}+\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}.$
We isolate $\eta_{k}$:
$\eta_{k}(1-\gamma^{T})\leq\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\Rightarrow\eta_{k}\leq\frac{1}{1-\gamma^{T}}\left(\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\right).$
We want to bound:
$\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}};$
using triangle inequality:
$\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}}\leq\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{\infty}\leq\eta_{K}+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}.$
Let us now plug in the bound for $\eta_{K}$:
$\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}}\leq\frac{1}{1-\gamma^{T}}\left(\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\right)+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}.$
The distribution mismatch is defined by $C_{k}:=\left\|\frac{d_{K}}{d_{k}}\right\|_{\infty}$. If $d_{k}$ is constructed by curriculum to smoothly interpolate toward $d_{K}$, and both $d_{K}$ and $d_{k}$ are supported on the same or growing state space. In our analysis, we use:
$\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{d_{K}}\leq C_{k}\cdot\|Q^{*}_{k}-Q^{\pi_{k}}_{k}\|_{\infty},$
where $\|\cdot\|_{d_{K}}$ is the $L_{2}$ norm under the distribution $d_{K}$ and $\|\cdot\|_{\infty}$ is the sup norm (worst-case).
In general, for any probability distribution $\mu$, and any function $f$,
[…]
So, in fact, we have the reverse:
$\|f\|_{d_{K}}\leq\|f\|_{\infty}\quad\Rightarrow\quad C_{k}\leq 1.$
If we do want to write
$\|f\|_{d_{K}}\leq C_{k}\cdot\|f\|_{\infty},$
then the tightest possible value for $C_{k}$ is exactly 1, and any value $C_{k}<1$ may be possible depending on the support of $f$ under $d_{K}$. This holds when later curriculum stages subsume earlier ones, a design principle in CRL. Hence, using $C_{k}\leq 1$ simplifies bounds without loosening them unnecessarily.
Because curriculum proceeds sequentially, and assuming that earlier tasks are easier and learned more accurately, we can sum the per-step bounds for curriculum steps:
$\mathcal{E}_{K}=\|Q^{*}_{K}-Q^{\pi_{K}}_{K}\|_{d_{K}}\leq\sum_{k=1}^{K}\left[\frac{1}{1-\gamma^{T}}\left(\frac{2\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\right)\right]+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}.$
This completes the proof.
Optionally, assuming $\gamma^{T}$ is small for large $T$, $1/(1-\gamma^{T})\leq 2$ leads to:
$\mathcal{E}_{K}\lesssim\sum_{k=1}^{K}\left(\frac{4\gamma(1-\gamma^{T})}{(1-\gamma)^{2}}\delta_{k}+\frac{4\gamma}{\beta(1-\gamma)^{2}}\right)+\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}.$
A.1.1 Theorem A.1 and Proof
Theorem A.1 (Finite-Sample Guarantee).
Consider the sequence of tasks $\{\mathcal{T}_{k}\}_{k=1}^{K}$, where the final task $\mathcal{T}_{K}$ has optimal action-value function $Q^{*}_{K}$. Suppose that $\pi_{K}$ is the final policy produced by CRL and assumptions hold for all curriculum stages. Following chen2022approximate , let $\mathcal{K}_{\mathcal{S}\mathcal{A}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|\times|\mathcal{S}||\mathcal{A}|}$ be a diagonal matrix and let $\mathcal{K}_{\mathcal{S}\mathcal{A},\min}$ be the minimal diagonal entry. $J_{k}$ is the number of critic updates per policy iteration at curriculum $k$. The bootstrapping parameter $n$ is chosen such that $\gamma_{c}:=\gamma^{n}/\sqrt{\mathcal{K}_{\mathcal{S}\mathcal{A},\min}}<1$. Then, when using constant stepsize $\alpha$ satisfying $\alpha(t_{\alpha}+n+1)\leq\frac{(1-\gamma_{c})\lambda_{\min}}{130L^{2}},$ we have for all $J_{k}\geq t_{\alpha}+n+1$:
$\mathbb{E}\left[\|Q^{*}_{K}-Q^{\pi_{K}}\|_{\infty}\right]$ $\leq\underbrace{\gamma^{TK}\|Q^{*}_{K}-Q^{\pi_{0}}\|_{\infty}}_{N_{1}}+\sum_{k=1}^{K}(\underbrace{\frac{2\gamma\mathcal{E}_{\text{approx},k}}{(1-\gamma)^{2}}}_{N_{2,1}}+\underbrace{\frac{2\gamma^{2}\mathcal{E}_{\text{bias},k}}{(1-\gamma)^{4}}}_{N_{2,2}})$
$+\underbrace{\sum_{k=1}^{K}\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}}_{N_{2,3}}+\underbrace{\sum_{k=1}^{K}\frac{70L_{k}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{\min}(1-\gamma_{c})(1-\gamma)^{3}}}_{N_{2,4}}$
$+\underbrace{\frac{2\gamma\beta K}{(1-\gamma)^{2}}}_{N_{3}}+\underbrace{\sum_{k=1}^{K-1}\|Q^{*}_{K}-Q^{*}_{k}\|_{d_{K}}}_{\text{Curriculumdiscrepancy}},$
where $c(\cdot,\cdot)$ and $\rho(\cdot,\cdot)$ are generalized importance sampling factors, $L_{k}$ is the parameter defined by $1+(\gamma\rho_{\text{max},k})^{n}$, $\mathcal{E}_{\text{approx},k}:=\sup_{\pi}\left\|Q^{\pi_{k}}_{c,\rho}-\Phi w^{\pi_{k}}_{c,\rho}\right\|_{\infty}$ is the critic’s function approximation error, $\mathcal{E}_{\text{bias},k}:=\max_{0\leq t\leq T}\max_{s\in\mathcal{S}}\left(1-\lambda(s)\right)\left\|\pi_{k,t}(\cdot|s)-\pi_{k,b}(\cdot|s)\right\|_{1}$ is the importance sampling bias at curriculum $k$, $\|Q^{*}_{K}-Q^{*}_{k}\|_{\infty}$ is the gap between the curriculum subtask and the final task, and $\lambda_{\min}$ is the mininum eigenvalue of the positive definite matrix $\Phi_{T}\mathcal{K}_{\mathcal{S}\mathcal{A}}\Phi$.
The terms $N_{1}$ and $N_{3}$ are the same as appeared in Thm. 3.1, and together capture the error in actor update. Term $N_{2,1}$ is the function approximation error and equals zero when using a complete basis. A term of similar form is present in all existing work studying RL with function approximation (munos2003error, ; agarwal2021theory, ). Term $N_{2,2}$ is the bias introduced by generalized importance sampling factors $c(\cdot,\cdot)$ and $\rho(\cdot,\cdot)$, and $N_{2,2}=0$ when $c(s,a)=\rho(s,a)=\pi(a|s)/\pi_{b}(a|s)$. Term $N_{2,3}$ represents the convergence bias in the critic and goes to zero geometrically fast as the inner loop iteration index $J_{k}$ goes to infinity. The term $N_{2,4}$ represents the variance in the critic and is proportional to $\sqrt{\alpha t_{\alpha}}=\mathcal{O}(\sqrt{\alpha\log(1/\alpha)})$, thus arbitrarily small given small enough stepsize $\alpha$. Together, $\{N_{2,i}\}_{i=1\sim 4}$ correspond to the second term in Thm. 3.1. Finally, the curriculum approximation error is the same as the last term in in Theorem 3.1. This result suggests that smoother curriculum trajectories and smaller approximation errors under finite samples, which is closely related with the stepsize and number of updates per iteration, can lead to better final policy performance.
Proof.
We proceed by mathematical induction across the curriculum steps. Let $\pi_{k,T}$ denote the final policy after $T$ iterations at curriculum step $k$. Base case: For the first curriculum step ($k=1$), we directly apply Theorem 4.1 from chen2022approximate :
$\mathbb{E}[\|Q_{1}-Q^{\pi_{1,T}}\|_{\infty}]$ $\leq\gamma^{T}\|Q_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$ (3)
$+\frac{2\gamma E_{approx,1}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,1}}{(1-\gamma)^{4}}$ (4)
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{1}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$ (5)
$+\frac{70L_{1}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}$ (6)
Inductive step: Assume that for curriculum step $k$, we have:
$\mathbb{E}[\|Q_{k}-Q^{\pi_{k,T}}\|_{\infty}]$ $\leq\gamma^{Tk}\|Q_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$
$+\sum_{j=1}^{k}\Bigg{(}\frac{2\gamma E_{approx,j}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,j}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{j}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{j}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}\Bigg{)}$
$+\sum_{j=1}^{k-1}\|Q_{k}-Q_{j}\|_{d_{k}}$
For curriculum step $k+1$, we initialize with policy $\pi_{k,T}$ and apply Theorem 4.1 from chen2022approximate :
$\mathbb{E}[\|Q_{k+1}-Q^{\pi_{k+1,T}}\|_{\infty}]$ $\leq\gamma^{T}\|Q_{k+1}-Q^{\pi_{k,T}}_{k+1}\|_{\infty}$
$+\frac{2\gamma E_{approx,k+1}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,k+1}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k+1}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{k+1}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}$
We need to relate $\|Q^{*}_{k+1}-Q^{\pi_{k,T}}_{k+1}\|_{\infty}$ to our induction hypothesis. By triangle inequality:
$\|Q_{k+1}-Q^{\pi_{k,T}}_{k+1}\|_{\infty}$ $\leq\|Q_{k+1}-Q_{k}\|_{\infty}+\|Q_{k}-Q^{\pi_{k,T}}_{k}\|_{\infty}+\|Q^{\pi_{k,T}}_{k}-Q^{\pi_{k,T}}_{k+1}\|_{\infty}$
$\leq\|Q_{k+1}-Q_{k}\|_{\infty}+\|Q^{*}_{k}-Q^{\pi_{k,T}}_{k}\|{\infty}+\frac{\gamma}{1-\gamma}\|r_{k}-r_{k+1}\|_{\infty}$
where the last inequality follows from the performance difference lemma with respect to rewards. For curriculum learning, we design the reward functions to satisfy $\|r_{k}-r_{k+1}\|_{\infty}\leq\delta_{r}$ for some small $\delta_{r}$. Thus:
$\|Q_{k+1}-Q^{\pi_{k,T}}_{k+1}\|_{\infty}$ $\leq\|Q_{k+1}-Q_{k}\|_{\infty}+\|Q_{k}-Q^{\pi_{k,T}}_{k}\|_{\infty}+\frac{\gamma\delta_{r}}{1-\gamma}$
Substituting our induction hypothesis:
$\mathbb{E}[\|Q_{k+1}-Q^{\pi_{k+1,T}}\|_{\infty}]$ $\leq\gamma^{T}\mathbb{E}[\|Q_{k+1}-Q_{k}\|_{\infty}+\|Q_{k}-Q^{\pi_{k,T}}_{k}\|_{\infty}+\frac{\gamma\delta_{r}}{1-\gamma}]$
$+\frac{2\gamma E_{approx,k+1}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,k+1}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k+1}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{k+1}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma}{\beta(1-\gamma)^{2}}$
$\leq\gamma^{T}\|Q_{k+1}-Q_{k}\|_{\infty}+\gamma^{T}\mathbb{E}[\|Q^{*}_{k}-Q^{\pi_{k,T}}_{k}\|_{\infty}]+\frac{\gamma^{T+1}\delta_{r}}{1-\gamma}$
$+\frac{2\gamma E_{approx,k+1}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,k+1}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k+1}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{k+1}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}$
Applying the induction hypothesis:
$\mathbb{E}[\|Q_{k+1}-Q^{\pi_{k+1,T}}\|_{\infty}]$ $\leq\gamma^{T}\|Q_{k+1}-Q_{k}\|_{\infty}+\gamma^{T}\cdot\gamma^{Tk}\|Q_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$
$+\gamma^{T}\sum_{j=1}^{k}\Bigg{(}\frac{2\gamma E_{approx,j}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,j}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{j}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{j}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma}{\beta(1-\gamma)^{2}}\Bigg{)}$
$+\gamma^{T}\sum_{j=1}^{k-1}\|Q_{k}-Q_{j}\|_{d_{k}}+\frac{\gamma^{T+1}\delta_{r}}{1-\gamma}$
$+\frac{2\gamma E_{approx,k+1}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,k+1}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k+1}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{k+1}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}$
Simplifying:
$\mathbb{E}[\|Q_{k+1}-Q^{\pi_{k+1,T}}\|_{\infty}]$ $\leq\gamma^{T(k+1)}\|Q_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$
$+\sum_{j=1}^{k+1}\Bigg{(}\frac{2\gamma E_{approx,j}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,j}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{j}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{j}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}\Bigg{)}$
$+\sum_{j=1}^{k}\|Q_{k+1}-Q_{j}\|_{d_{k+1}}$
where we used the fact that
$\|Q^{*}_{k+1}-Q^{*}_{k}\|_{d_{k+1}}+\gamma^{T}\sum_{j=1}^{k-1}\|Q^{*}_{k}-Q^{*}_{j}\|_{d_{k}}\leq\sum_{j=1}^{k}\|Q^{*}_{k+1}-Q^{*}_{j}\|_{d_{k+1}}$ (7)
due to the triangle inequality and the curriculum design. By induction, for the final curriculum step $K$, we have:
$\mathbb{E}[\|Q_{K}-Q^{\pi_{K}}\|_{\infty}]$ $\leq\gamma^{TK}\|Q_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$
$+\sum_{k=1}^{K}\Bigg{(}\frac{2\gamma E_{approx,k}}{(1-\gamma)^{2}}+\frac{2\gamma^{2}E_{bias,k}}{(1-\gamma)^{4}}$
$+\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$
$+\frac{70L_{k}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}+\frac{2\gamma\beta}{(1-\gamma)^{2}}\Bigg{)}$
$+\sum_{k=1}^{K-1}\|Q_{K}-Q_{k}\|_{d_{K}}$
This completes the proof.
A.1.2 Proof of Theorem 3.2
Proof.
From the Finite-Sample Theorem for CRL, we aim to control all error terms to achieve the desired accuracy. We allocate the total error budget $\epsilon$ across the $K$ curriculum steps, with $\epsilon_{k}$ being the error allocation for step $k$, such that $\sum_{k=1}^{K}\epsilon_{k}\leq\epsilon$. For each curriculum $k$, we need to control the following terms:
$\gamma^{TK}\|Q^{*}_{1}-Q^{\pi_{0}}_{1}\|_{\infty}$ $\leq\frac{\epsilon}{4}$ (8)
$\frac{6(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k}-(t_{\alpha}+n+1)]}}{(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}$ $\leq\frac{\epsilon_{k}}{4}$ (9)
$\frac{70L_{k}[\alpha(t_{\alpha}+n+1)]^{1/2}}{\lambda_{min}(1-\gamma_{c})(1-\gamma)^{3}}$ $\leq\frac{\epsilon_{k}}{4}$ (10)
$\frac{2\gamma\beta}{(1-\gamma)^{2}}$ $\leq\frac{\epsilon_{k}}{4}$ (11)
We can solve each of these constraints. For the first constraint, we need:
$T\geq\frac{1}{K}\log_{\gamma}\left(\frac{\epsilon}{4\|Q^{*}_{1}-Q^{\pi_{0}}_{1}\|_{\infty}}\right)$
Since $\|Q^{*}_{1}-Q^{\pi_{0}}_{1}\|_{\infty}\leq\frac{1}{1-\gamma}$, we have:
$T\geq\frac{1}{K}\log\gamma\left(\frac{\epsilon(1-\gamma)}{4}\right)=O\left(\frac{\log(1/\epsilon)}{K}\right)$
For the second constraint, we need:
$(1-(1-\gamma_{c})\lambda_{min}\alpha)^{\frac{1}{2}[J_{k}-(t_{\alpha}+n+1)]}$ $\leq\frac{\epsilon_{k}(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}{24},$
thus by taking log,
$\frac{1}{2}[J_{k}-(t_{\alpha}+n+1)]\log(1-(1-\gamma_{c})\lambda_{min}\alpha)$ $\leq\log\left(\frac{\epsilon_{k}(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}{24}\right),$
thus
$J_{k}$ $\geq(t_{\alpha}+n+1)+\frac{2\log\left(\frac{\epsilon_{k}(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}{24}\right)}{\log(1-(1-\gamma_{c})\lambda_{min}\alpha)}.$
Using $\log(1-x)\approx-x$ for small $x$, and $\alpha(t_{\alpha}+n+1)\leq\frac{(1-\gamma_{c})\lambda_{min}}{130L_{k}^{2}}$, we have:
$J_{k}$ $\geq(t_{\alpha}+n+1)-\frac{2\log\left(\frac{24}{\epsilon_{k}(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}\right)}{(1-\gamma_{c})\lambda_{min}\alpha}$
$=(t_{\alpha}+n+1)+\frac{2\log\left(\frac{\epsilon_{k}(1-\gamma)^{3}(1-\gamma_{c})^{1/2}\lambda_{min}^{1/2}}{24}\right)}{(1-\gamma_{c})\lambda_{min}\alpha}$
For the third constraint, we need:
$\alpha(t_{\alpha}+n+1)$ $\leq\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{4900L_{k}^{2}}$
$=O\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}}\right)$
For the fourth constraint, we need:
$\beta\leq\frac{\epsilon_{k}(1-\gamma)^{2}}{8\gamma}$
Combining these constraints, the dominant factor in the sample complexity comes from the third constraint, which gives us:
$\alpha\leq O\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}(t_{\alpha}+n+1)}\right)$
The mixing time $t_{\alpha}$ is $O(\log(1/\alpha))$, which gives us:
$\alpha$ $\leq O\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}(\log(1/\alpha)+n+1)}\right)$
This implies:
$\alpha$ $=O\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}\log(1/\alpha)}\right)$
$\alpha\log(1/\alpha)$ $=O\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}}\right)$
Using the Lambert W function, we can solve for $\alpha$:
$\alpha$ $=\Theta\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}\log\left(\frac{L_{k}^{2}}{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}\right)}\right)$
$=\tilde{\Theta}\left(\frac{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}{L_{k}^{2}}\right)$
where $\tilde{\Theta}$ hides logarithmic factors. The number of samples required for curriculum step $k$ is:
$M_{k}$ $=J_{k}\cdot T$
$=O\left((t_{\alpha}+n+1)+\frac{\log(1/\epsilon_{k})}{(1-\gamma_{c})\lambda_{min}\alpha}\right)\cdot O\left(\frac{\log(1/\epsilon)}{K}\right)$
$=O\left(\frac{\log(1/\alpha)+n+1}{K}\log(1/\epsilon)+\frac{\log(1/\epsilon_{k})\log(1/\epsilon)}{K(1-\gamma_{c})\lambda_{min}\alpha}\right)$
$=O\Bigg{(}\frac{\log(1/\epsilon)\log\left(\frac{L_{k}^{2}}{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}\right)+n\log(1/\epsilon)}{K}$
$+\frac{\log(1/\epsilon_{k})\log(1/\epsilon)L_{k}^{2}\log\left(\frac{L_{k}^{2}}{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{2}\lambda_{min}^{2}}\right)}{K\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\Bigg{)}$
$=\tilde{O}\left(\frac{n\log(1/\epsilon)}{K}+\frac{\log(1/\epsilon_{k})\log(1/\epsilon)L_{k}^{2}}{K\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)$
The second term dominates, giving us:
$M_{k}=\tilde{O}\left(\frac{\log^{2}(1/\epsilon_{k})\log(1/\epsilon)L_{k}^{2}}{K\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)$ (12)
The total number of samples across all curriculum steps is:
$M_{CRL}$ $=\sum_{k=1}^{K}M_{k}$
$=\sum_{k=1}^{K}\tilde{O}\left(\frac{\log^{2}(1/\epsilon_{k})\log(1/\epsilon)L_{k}^{2}}{K\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)$
$=\tilde{O}\Bigg{(}\frac{\log(1/\epsilon)}{K}\sum_{k=1}^{K}\frac{\log^{2}(1/\epsilon_{k})L_{k}^{2}}{\epsilon_{k}^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\Bigg{)}$
Assuming $\epsilon_{k}=\frac{\epsilon}{K}$ for all $k$ (uniform error allocation), we get:
$M_{CRL}$ $=\tilde{O}\left(\frac{\log(1/\epsilon)}{K}\sum_{k=1}^{K}\frac{\log^{2}(K/\epsilon)L_{k}^{2}}{(\epsilon/K)^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)$
$=\tilde{O}\left(\frac{\log(1/\epsilon)\log^{2}(K/\epsilon)K}{\epsilon^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\sum_{k=1}^{K}L_{k}^{2}\right)$
$=\tilde{O}\left(\frac{\log^{3}(1/\epsilon)K}{\epsilon^{2}(1-\gamma)^{6}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\sum_{k=1}^{K}L_{k}^{2}\right)$
Adding the dependence on the bootstrapping parameter $n$, we have:
$M_{CRL}=O\left(\sum_{k=1}^{K}\frac{\log^{3}(1/\epsilon_{k})}{\epsilon_{k}^{2}}\cdot\tilde{O}\left(\frac{L_{k}^{2}n}{(1-\gamma)^{7}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right)\right)$ (13)
This completes the proof for the first half of the Theorem. Now we move on to prove the second half of the Theorem.
The sample complexity for the final task $K$ is:
$M_{K}=O\left(\frac{\log^{3}(1/\epsilon_{K})}{\epsilon_{K}^{2}}\cdot L_{K}^{2}\cdot\frac{C}{(1-\gamma)^{7}(1-\gamma_{c})^{3}\lambda_{min}^{3}}\right).$
We can factor this term out of the total sum:
$M_{CRL}=M_{K}\cdot\Bigg{(}1+\sum_{k=1}^{K-1}\frac{L_{k}^{2}\cdot\log^{3}(1/\epsilon_{k})\cdot\epsilon_{K}^{2}}{L_{K}^{2}\cdot\log^{3}(1/\epsilon_{K})\cdot\epsilon_{k}^{2}}\Bigg{)}.$
Define the curriculum efficiency factor:
$CEF=1+\sum_{k=1}^{K-1}\frac{L_{k}^{2}\cdot\log^{3}(1/\epsilon_{k})\cdot\epsilon_{K}^{2}}{L_{K}^{2}\cdot\log^{3}(1/\epsilon_{K})\cdot\epsilon_{k}^{2}}.$
This factor represents the ratio of total curriculum sample complexity to the sample complexity of just the final task. We make two structure assumptions about our curriculum design, geometric error allocation and $L_{k}$ progression, $\epsilon_{k}=\epsilon_{K}\cdot e^{K-k}$ and $L_{k}=L_{K}/l^{K-k}$, where $e>1$. This reflect a curriculum that gradually increases in difficulty while allowing larger errors in earlier stages. For a well-designed curriculum where early tasks are simpler than later ones, $l>1$.
For the error ratio term:
$\frac{\epsilon_{K}^{2}}{\epsilon_{k}^{2}}=\frac{\epsilon_{K}^{2}}{(\epsilon_{K}\cdot e^{K-k})^{2}}=e^{-2(K-k)},$
and for the $L_{k}$ ratio term:
$\frac{L_{k}^{2}}{L_{K}^{2}}=\frac{L_{k}^{2}}{(L_{k}\cdot l^{K-k})^{2}}=l^{-2(K-k)}.$
For the logarithmic term:
$\log(1/\epsilon_{k})=\log(1/(\epsilon_{K}\cdot e^{K-k}))=\log(1/\epsilon_{K})-(K-k)\log(e).$
Substituting these expressions into $CEF$:
$CEF=1+\sum_{k=1}^{K-1}\frac{l^{-2(K-k)}\cdot[\log(1/\epsilon_{K})-(K-k)\log(e)]^{3}\cdot e^{-2(K-k)}}{\log^{3}(1/\epsilon_{K})}.$
Since $e>1$ and $K\geq k$, $(K-k)\log(e)\geq 0$. Thus we have
$CEF$ $=1+\sum_{k=1}^{K-1}\frac{[\log(1/\epsilon_{K})-(K-k)\log(e)]^{3}}{\log^{3}(1/\epsilon_{K})}\cdot(el)^{-2(K-k)}$
$\leq 1+\sum_{k=1}^{K-1}(el)^{-2(K-k)}.$
The sum is a geometric series:
$\sum_{k=1}^{K-1}(el)^{-2(K-k)}=\sum_{k=1}^{K-1}(el)^{2(k-K)}=(el)^{2(1-K)}\frac{1-(el)^{2(K-1)}}{1-(el)^{2}}=\frac{(el)^{2(1-K)}-1}{1-(el)^{2}}$
CRL is more sample-efficient than direct learning when:
$M_{CRL}<M_{Direct}$
Since $M_{CRL}=M_{K}\cdot CEF$ and for a comparable direct learning approach, $M_{Direct}=M_{K}\cdot m$ where $m$ represents the relative difficulty factor of direct learning, our condition becomes:
$CEF<m.$
Thus for $M_{CRL}<M_{Direct}$, substituting expressions it becomes:
$M_{CRL}=M_{K}\cdot CEF\leq M_{K}\cdot(1+\frac{(el)^{2(1-K)}-1}{1-(el)^{2}})<M_{K}\cdot m=M_{Direct}.$
Thus
$M_{\text{CRL}}<M_{\text{Direct}}\iff\frac{(el)^{2(1-K)}-1}{1-(el)^{2}}<m-1.$
This completes the proof.
Appendix B Dataset Details
In this section we provide details of the datasets used for evaluation. We categorize the datasets into two categories, datasets that contain human annotated difficulties and others that do not.
B.1 Datasets with Human Annotated Difficulties
Blocksworld (valmeekam2023planning, ) is a dataset used to evaluate the planning capabilities of LLMs. Each task involves transitioning from an initial block configuration to a target configuration, which requires LLMs to generate a sequence of actions, or plan to achieve the goal. Tasks become more difficult as the required number of steps increases, since the model must reason over longer sequences and maintain correct intermediate states. To study this, we group tasks into four in-distribution difficulty levels: Trivial with 1 step, Easy with 2 steps, Medium with 4 steps, and Hard with 6 steps. Additionally, we include an out-of-distribution (OOD) split with 8-step plans to test generalization beyond the training distribution. Trivial tasks are especially simple because the model only needs to predict one correct action out of four possible choices to complete the plan. This setup allows the LLM to first grasp fundamental planning mechanics, which can then be leveraged to learn more complex multi-step tasks.
Countdown (gandhi2024stream, ) is a task where the model must reach a target value by combining given numbers using basic arithmetic operations. While the original dataset uses four numbers per instance, we extend it to create a range of difficulty levels based on the number of input numbers, mainly, Trivial (2), Easy (3), Medium (4), Hard (5), and OOD (6). As the number of inputs increases, the space of possible operation sequences grows rapidly, making it harder for the model to identify the correct combination and order of operations. In contrast, the trivial setting is extremely simple, requiring just one operation between two numbers to reach the target, allowing the model to first learn basic arithmetic before scaling to more complex multi-step problems.
MATH (hendrycksmath2021, ) is a benchmark of 7,500 training and 5,000 test problems covering high-school level mathematics, with each problem labeled from Level 1 (easiest) to Level 5 (hardest). The dataset covers topics such as algebra, geometry, number theory, and probability, including step-by-step solutions. We create a difficulty-based setup using the existing labels, specifically, Trivial (Level 1), Easy (Level 2), Medium (Level 3), Hard (Level 4), and OOD (Level 5). As difficulty increases, problems require more complex reasoning, multi-step solutions, and deeper mathematical understanding. Trivial problems are typically short and rely on basic techniques, making them ideal for teaching foundational reasoning. We use Levels 1 through 4 for training and treat Level 5 as out-of-distribution to assess generalization to the most difficult problems.
We provide the summery of the difficulty splits for each dataset in Table 5.
B.2 Datasets without Human Annotated Difficulties
GSM8K (cobbe2021gsm8k, ) is a dataset of high-quality, linguistically diverse grade school math word problems designed to evaluate multi-step arithmetic reasoning. Each problem typically requires between two and eight steps involving basic arithmetic operations such as addition, subtraction, multiplication, and division. To assess performance across varying difficulty levels, we create a 4-way split into Trivial, Easy, Medium, and Hard, based on model error rates, as described in Section. 4 and illustrated in Figure. 4.2
AQuA (ling2017program, ) is a dataset of algebraic word problems with multiple-choice answers and detailed rationales, designed to test arithmetic reasoning and symbolic manipulation. For our experiments, we randomly sample 5,000 problems for training. Similar to GSM8K, we use model error rates to define four difficulty levels, Trivial, Easy, Medium, and Hard (see Figure. 6).
Since both GSM8K and AQuA lack explicit difficulty annotations, we construct difficulty splits using model error rates and do not include an out-of-distribution (OOD) category for these datasets.
Table 5: Difficulty splits for datasets with human-annotated difficulty levels. Each dataset is categorized based on task-specific properties, specifically, plan length for Blocksworld, number of operands for Countdown, and problem level for MATH.
Difficulty Blocksworld Countdown MATH
(Plan Length) (Num. Operands) (Problem Level)
Trivial (T) 1 2 1
Easy (E) 2 3 2
Medium (M) 4 4 3
Hard (H) 6 5 4
OOD 8 6 5
Figure 6: AQuA Difficulty distribution based on error rates, grouped into quartiles.
Appendix C Model License Information
For our experiments, we use LLaMA 3.2 3B Instruct, Qwen 2.5 1.5B Instruct, and Qwen 2.5 3B Instruct. Qwen 2.5 1.5B Instruct is released under the Apache 2.0 License, Qwen 2.5 3B Instruct under the Qwen Research License, and LLaMA 3.2 3B Instruct under the Meta Community License.
Appendix D Implementation Details
In this section, we provide implementation details for E2H Reasoner. We begin by describing the hardware setup, followed by training parameters, and finally the inference settings used to report the final results.
D.1 Hardware Details
For our experiments, we use up to three 80GB NVIDIA A100 GPUs, particularly for the 3B parameter models. One GPU is dedicated to vLLM for fast inference, while the remaining two are used for model training. On average our experiments require anywhere between 16-18 hours for training.
D.2 Training Details
We use GRPO as the policy optimization algorithm, combined with parameter-efficient fine-tuning using LoRA. All models are trained for up to 1600 GRPO steps, with hyperparameters listed in Table 6. For fair comparison, the same configuration is used in our SFT experiments. All training is conducted using the TRL library from Hugging Face.
Table 6: Training hyperparameters used for GRPO post-training and LoRA adaptation.
Component Parameter Value
GRPO learning_rate 1e-6
lr_scheduler_type cosine
per_device_train_batch_size 2
gradient_accumulation_steps 4
gradient_checkpointing true
max_steps 1600
bf16 true
tf32 true
num_generations 8
beta 0.001
use_vllm true
vllm_gpu_memory_utilization 0.2
LoRA r 32
alpha 64
dropout 0.1
target_modules q_proj, v_proj
task_type CAUSAL_LM
D.3 Inference Details
For inference, we use a temperature of 0.0 in all experiments to ensure deterministic outputs and reproducibility. For pass@$k$ evaluations, we enable sampling with a temperature of 0.7, top_p of 0.9, and top_k of 50, where top_k controls the number of candidate tokens considered at each decoding step.
Appendix E Comparisons with SFT
We compare vanilla supervised fine-tuning (SFT) with RL-based post-training methods (see Table 7). The results show that SFT performance varies significantly across tasks. For instance, in Blocksworld (valmeekam2023planning, ), where most problems involve fewer than four blocks, SFT performs well—the LLM learns effectively due to the small gap between the training and test distributions. This aligns with prior work suggesting that SFT is most effective when the training data is high quality and closely aligned with the downstream task (muennighoff2025s1, ; chu2025sft, ). In contrast, on tasks like Countdown, where the problem distribution is much broader, SFT performs notably worse than RL-based post-training. This gap is especially clear in the out-of-distribution (OOD) setting, where SFT fails to solve any Countdown OOD examples, unlike in Blocksworld, where its performance generalizes more successfully. We highlight these two tasks to illustrate how SFT performance can vary across distributions, in contrast to the more consistent behavior of RL-based post-training methods.
Table 7: Comparison of RL-based post-training and supervised fine-tuning (SFT). SFT performance varies noticeably across tasks, highlighting its inconsistency in generalizing across domains.
Models Methods Blocksworld Countdown
Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD
Qwen 2.5 / 1.5B Instruct SFT 100.0 97.8 88.1 55.3 16.5 97.4 41.8 14.2 4.6 0.0
GRPO 98.0 100.0 84.5 26.3 5.3 96.1 64.9 28.8 18.1 9.2
E2H-G 98.0 100.0 95.3 32.9 7.3 95.2 84.1 48.1 28.1 14.2
E2H-C 100.0 100.0 15.5 0.0 0.0 96.7 64.0 25.9 15.8 6.4
Qwen 2.5 / 3B Instruct SFT 100.0 100.0 94.0 62.5 27.8 99.9 63.9 30.8 13.3 0.0
GRPO 100.0 100.0 88.1 38.8 14.6 99.6 89.3 73.4 41.7 9.4
E2H-G 100.0 100.0 96.4 53.3 23.2 98.5 90.8 71.0 43.5 19.4
E2H-C 100.0 100.0 94.1 52.0 22.5 100.0 90.4 69.7 35.7 12.6
LLaMA 3.2 / 3B Instruct SFT 100.0 97.8 96.4 72.3 37.8 100.0 66.6 27.2 12.1 0.0
GRPO 100.0 100.0 94.1 38.9 13.3 99.9 89.5 71.6 47.9 2.7
E2H-G 100.0 100.0 98.8 44.1 17.6 95.0 89.9 73.3 46.5 24.3
E2H-C 100.0 0.0 0.0 0.0 0.0 100.0 55.3 0.0 0.0 0.0
Appendix F Additional Results and Analysis of Scheduling
Table 8: Effect of scheduling strategies in LLM post-training. We compare balanced scheduling, traditional curriculum learning (CL), and our proposed E2H Reasoner variants, namely, E2H-G and E2H-C. CoT is reported as a reference.
Blocksworld Countdown MATH
Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD Trivial Easy Med Hard OOD
Qwen 2.5 3B Instruct
CoT 1.0 6.7 7.1 1.0 0.0 24.0 15.9 4.4 0.5 0.0 68.0 55.8 48.8 33.4 20.0
Balanced 100.0 100.0 88.1 38.8 14.6 99.6 86.9 57.8 23.6 7.8 90.1 78.5 71.9 57.4 36.8
CL 100.0 100.0 77.3 42.1 20.5 39.0 91.0 72.9 38.7 12.5 90.6 78.9 72.5 57.2 36.6
E2H-G (0.25, 0.75) 100.0 100.0 96.4 53.3 23.2 98.5 90.8 71.0 43.5 19.4 90.4 81.0 71.5 58.2 37.2
E2H-G (0.5, 0.5) 100.0 100.0 89.2 32.9 10.6 99.6 89.3 73.4 41.7 9.4 91.1 78.2 70.5 56.6 35.0
E2H-G (0.75, 0.25) 100.0 100.0 42.9 3.9 2.0 100.0 87.8 52.3 16.7 6.5 90.0 79.3 70.6 57.1 35.9
E2H-C 100.0 100.0 94.1 52.0 22.5 100.0 90.4 69.7 35.7 12.6 90.4 80.0 73.4 58.9 38.1
LLaMA 3.2 3B Instruct
CoT 24.0 0.0 1.2 1.0 0.0 37.1 4.6 0.3 0.0 0.0 65.9 44.6 35.2 24.1 13.6
Balanced 100.0 100.0 94.1 38.9 13.3 99.9 89.5 71.6 47.9 2.7 65.9 47.0 36.0 22.0 10.2
CL 100.0 0.0 0.0 0.0 0.0 17.2 36.0 22.7 11.2 4.1 74.1 54.1 43.9 28.0 12.5
E2H-G (0.25, 0.75) 100.0 100.0 98.8 44.1 17.6 95.0 89.9 73.3 46.5 24.3 78.7 58.4 46.4 32.3 14.5
E2H-G (0.5, 0.5) 100.0 100.0 88.1 26.3 9.6 94.4 86.7 65.1 27.9 2.3 66.1 45.0 36.2 24.4 11.4
E2H-G (0.75, 0.25) 100.0 97.8 75.0 21.1 4.0 98.9 89.4 48.6 0.7 0.0 63.8 48.2 35.6 24.0 8.6
E2H-C 100.0 100.0 15.5 0.0 0.0 96.7 64.0 25.9 15.8 6.4 84.6 69.6 63.0 47.6 28.6
In this section, we expand on the results presented in the main paper by reporting all three parameter settings of Gaussian scheduling in Table 8. Similarly, for Qwen 2.5 1.5B Instruct we include the results for all Gaussian scheduling variants on GSM8K and AQuA in Table 9.
Table 9: Expanded results for Qwen 2.5 1.5B Instruct showing all E2H-G (Gaussian scheduling) variants on GSM8K and AQuA. We provide a comparison against Balanced and E2H-C (Cosine scheduling).
AQuA GSM8K
Trivial Easy Med Hard Avg Trivial Easy Med Hard Avg
Balanced 95.8 68.0 48.6 21.0 63.3 99.0 95.3 84.1 49.9 77.1
E2H-C 86.1 72.2 48.6 26.3 62.5 98.0 95.3 83.9 46.6 75.7
E2H-G (0.25, 0.75) 93.0 70.8 40.2 23.6 61.4 98.5 96.3 83.6 51.2 77.6
E2H-G (0.5, 0.5) 90.2 81.9 43.0 34.2 66.1 97.6 94.7 89.0 51.8 78.7
E2H-G (0.75, 0.25) 83.3 79.1 54.5 26.3 64.1 98.0 97.3 85.9 47.7 77.1
Finally, we provide a qualitative analysis of scheduling techniques in Table 10, highlighting their respective strengths and weaknesses. The results show that different schedulers are suited to different tasks and model scales. Inspired by the effectiveness of E2H Reasoner, we hope this motivates the community to explore more sophisticated scheduling strategies for LLM post-training.
Table 10: Qualitative comparison of scheduling strategies, outlining their strengths and weaknesses in the context of LLM post-training.
Scheduling Strengths Weaknesses
Balanced
1. Default in LLM post-training.
2. Parameter-free and easy to use.
1. Assumes uniform difficulty across the dataset.
2. Often fails to improve reasoning on hard tasks.
Traditional
1. Simple and parameter-free.
2. Easy to implement.
1. Can cause overfitting to easy tasks.
2. May lead to forgetting earlier tasks.
E2H-C
1. Parameter-free and simple to apply.
2. Suitable for tasks with similar zero-shot
      performance across difficulty levels.
1. May overfit to easy tasks when rewards for hard
     tasks are sparse.
E2H-G
1. Effective across tasks and models.
2. Enables fine-grained control.
1. Requires tuning two hyperparameters.
Appendix G Societal Impacts
Our work introduced E2H Reasoner, a curriculum-based reinforcement learning (RL) approach designed to enhance the reasoning capabilities of small-scale large language models (LLMs). This advancement holds significant societal implications. Enhanced reasoning in LLMs can improve decision-making processes in critical domains such as healthcare, education, and legal systems, where nuanced understanding is paramount. Moreover, by empowering smaller models, E2H Reasoner promotes broader accessibility to advanced AI capabilities. Although our tasks focus on reasoning, future extensions involving real-world interaction could pose risks, including potential misuse of language models.